OpenStudy (sedatefrog712):

MEDAL A sandbag was thrown downward from a building. The function f(t) = -16t2 - 64t + 80 shows the height f(t), in feet, of the sandbag after t seconds: Part A: Factor the function f(t) and use the factors to interpret the meaning of the x-intercept of the function. (4 points) Part B: Complete the square of the expression for f(x) to determine the vertex of the graph of f(x). Would this be a maximum or minimum on the graph? (4 points) Part C: Use your answer in part B to determine the axis of symmetry for f(x)? (2 points)

OpenStudy (sedatefrog712):

@ganeshie8

OpenStudy (anonymous):

-16t^2+-64+80 A.) Factor -16(T^2+4x-5) -16(T+5)(T-1) x= (-8,1) B.) -16(T^2+4x+4)+64+40 -16(T+2)^2+104 Vertex=(-2,104) C.) X=-b/2a -(-64)/2(-16) Symmetry x=-2

OpenStudy (anonymous):

Sorry for the time, but let me know if you need any more clarification

OpenStudy (sedatefrog712):

@n648c788 thank you but i have a coupple more questions for part a and b

OpenStudy (anonymous):

Hey sorry I did not see you messages, I am still new to the site